Hi i need help with this question on logarithms.

Evaluate this logarithm.
log base 8 of 1/256
Thank you very much

Answer by dkspoet(13)

Here Log8 denotes Log to the base 8.
Log8 (1/256)
= Log8 (1/2)^8
= 8 Log8 (1/2)
= 8(Log8 1 - Log8 2)
= 8(0 - Log8 2)
= - 8/3 Log8 2^3
= - 8/3 Log8 8
= - 8/3